OpenStudy (anonymous):

Find the product of (x + 7)(x - 7). x2 - 14x + 49 x2 - 49 x2 + 14x + 49 x2 + 49

OpenStudy (gorv):

just expand it

OpenStudy (gorv):

x*(x-7)+7*(x-7) can u do it now ...just simple multiplication

OpenStudy (anonymous):

x^2 + 14x + 49?

OpenStudy (anonymous):

or would it just be x^2 + 49?

OpenStudy (anonymous):

@gorv

OpenStudy (skullpatrol):

$$\Huge (x + 7)(x - 7)=x^2 - 7^2$$

OpenStudy (anonymous):

so x^2 - 49

OpenStudy (skullpatrol):

$$\Huge (a+b)(a-b)=a^2-b^2$$

OpenStudy (skullpatrol):

Yes

OpenStudy (anonymous):

thanks both of you :D

OpenStudy (skullpatrol):

Thanks for asking :-)